OpenStudy (anonymous):

Just need to make sure. evaluate the derivative dy/dx at the point (0,0), for the equation 2x-5x^3y^2+4y=0

OpenStudy (anonymous):

@RBauer4

OpenStudy (anonymous):

i think it is undefined

OpenStudy (anonymous):

my options are 3, -1/2, undefinded, 6, and none of these

OpenStudy (jdoe0001):

hmm what did you get for the \(\cfrac{dy}{dx}\) of \(2x-5x^3y^2+4y=0\) ?

OpenStudy (anonymous):

I got 2-15x^2y^2

OpenStudy (jdoe0001):

2x - 5x^3y^2 + 4y ^ ^ ^ power rule product rule power rule

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf 2x-5x^3y^2+4y=0\qquad \cfrac{dy}{dx}\implies 2-(15x^2\cdot y^2+5x^3\cdot 2y)+4\\ \quad \\ \cfrac{dy}{dx}\implies -15x^2y^2-10x^3y+6\)

OpenStudy (anonymous):

oh ok i did it wrong :/

OpenStudy (anonymous):

so now I insert 0 in x and y?

OpenStudy (jdoe0001):

well... that's the derivative equation, the equation for the slope, now to get the slope itself at (0, 0), just set x =0 that'd yield a "y" of 6

OpenStudy (anonymous):

so the answer is 6? ... wow I did that very wrong

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thank you so much. :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

could you help me with one more?